OpenStudy (anonymous):

1. Determine the slope of the line that passes through the points (2, 2) and (4, -5).

OpenStudy (anonymous):

y2-y1/x2-x1= -5-2/4-2 =-7/2

myininaya (myininaya):

\[\frac{y_1-y_2}{x_1-x_2}\] plug in the following \[(x_1=2,y_1=2) \text{ and } (x_2=4,y_2=-5)\]

OpenStudy (anonymous):

to late i hav failed this one to

myininaya (myininaya):

Ok please stop saying that word. Obviously we have it censored for a reason. Thanks.